Circular No. 6484

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GALACTIC CENTER T. Strohmayer, Laboratory for High Energy Astrophysics and Univer-

sities Space Research Association (LHEA/USRA); U. Lee, Tohoku University, Sendai; and K. Jahoda, LHEA, Goddard Space Flight Center, report: “We have detected x-ray brightness oscillations with a frequency of 589 Hz during three type-I x-ray bursts from a burster in the galactic-center region. The bursts ocurred on May 15.814, June 4.612, and June 19.414 UT during routine monitoring observations of GRO J1744–28 by the Rossi X-ray Timing Explorer (RXTE). The oscillations have amplitudes (rms) in the range 2–4 percent (above 8 keV) and are detected only in the few seconds after burst peak. Each burst shows evidence for modest photospheric radius expansion. The current data constrain the source position to a 0^{o}.15-wide region whose shape is approximately that of one quarter of an annulus. The center line of the annular region extends from (α, δ) = (17^{h}43^{m}40^{s}.8, −29^{o}03′00′′, equinox 2000.0) to (17^{h}45^{m}00^{s}.0, −29^{o}16′48′′), with midpoint at (17^{h}44^{m}07^{s}.2, −29^{o}09′00′′), strongly excluding GRO J1744–28 as the burst source. The known burster nearest to this region is MXB 1743–29, but at present this association is only suggestive. This represents the third detection by RXTE of a near-millisecond oscillation from an x-ray bursting low-mass x-ray binary (see IAUC 6387, 6437).”

GRO J1744–28 T. Augusteijn and C. Lidman, European Southern Observatory; and

P. Blanco, University of California, San Diego; report: “A re-analysis of the Feb. 8 K-band image taken of the field of the transient GRO J1744–28 (IAUC 6321) shows that the source that was identified as the infrared counterpart of the transient (IAUC 6369) may not be real. The image is the superposition of ten individual, mutually-shifted images. The source was in the field-of-view of seven of these; however, it only appears in two. The detections do not coincide with any recorded bursts in the Ulysses GRB data (K. Hurley, private communication), and we suspect that the detections may be artifacts of the detector. Further analysis of the Feb. 8 and Mar. 28 images do not show any source in the vicinity of the x-ray position down to K ∼ 16 and varying by > 0.5 mag. We encourage continuous deep-infrared imaging to secure the identification of the counterpart of GRO J1744–28.”

1996 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green